# EchoAgent: Towards Reliable Echocardiography Interpretation with “Eyes”, “Hands” and “Minds”

### 07 Apr 2026 v1

@@ 2.1 Specialized task-specific networks for Echo
Developing DL networks to automate Echo analysis has become a trendy research objective [10, 18, 26, 42, 6]. Plenty of work has focused on specific sub-tasks, primarily including view identification and anatomical structure segmentation. Early approaches employed CNNs to identify views [21, 43], while recent advancements have utilized multi-instance learning[33] or pretrained foundation models (FMs) [9, 38] to capture complex visual patterns for accurate view prediction. For segmentation, U-Net remains the strong baseline [23], with attention mechanisms being adapted to improve boundary delineation for cardiac structures [15, 28, 31, 44]. Moving forward, the emergence of FMs have also spurred remarkable advances in Echo segmentation, where FM-based models such as MemSAM [10] and EchoONE [18] facilitate precise segmentation for Echo. While these models excel in their designated tasks, they operate with limited “eyes” and “hands”, requiring separate manual assistance for final diagnostic conclusion, which is not favorable for comprehensive Echo interpretation.
@@ 2.3 Towards domain-expertise agentic system
Recent days have witnessed growing interest in developing autonomous agentic systems, driven by the unparalleled tool-calling capability of MLLMs most recently [3, 32]. This promotes the development of agents to transcend the conversational nature of standard MLLMs by actively orchestrating external tools to acquire quantitative evidence and execute complex workflows [11, 30, 7]. For instance,  [11] introduces a pipeline MedRAX, integrating specialized CXR tools with MLLMs to address complex chest X-ray analysis. In the more generalized CT and X‑ray domains, [30] enhances diagnostic performance by explicitly infusing MLLMs with knowledge from domain‑specialized vision models. Meanwhile, [7] presents a collaborative‑learning framework that adaptively incorporates pre-trained single-modal medical large models with small multimodal models to boost multimodal diagnosis. Despite these advances, we still lack a specialized agent for reliable Echo interpretation, since Echo is a unique imaging modality, characterized by its inherent multi‑view interdependency, complex anatomical geometry, and ambiguous structure boundaries. These challenges necessitate a specialized synergy of Echo expertise with perceptual, operational, and reasoning capabilities, which existing agents are not designed to fulfill.

### 16 Apr 2026 v2

@@ 2.1 Specialized task-specific networks for Echo
Developing DL networks to automate Echo analysis has become a trendy research objective [12, 20, 28, 44, 2, 7]. Plenty of work has focused on specific sub-tasks, primarily including view identification and anatomical structure segmentation. Early approaches employed CNNs to identify views [23, 45], while recent advancements have utilized multi-instance learning[35] or pretrained foundation models (FMs) [10, 40] to capture complex visual patterns for accurate view prediction. For segmentation, U-Net remains the strong baseline [25], with attention mechanisms being adapted to improve boundary delineation for cardiac structures [17, 30, 33, 46]. Moving forward, the emergence of FMs have also spurred remarkable advances in Echo segmentation, where FM-based models such as MemSAM [12] and EchoONE [20] facilitate precise segmentation for Echo. While these models excel in their designated tasks, they operate with limited “eyes” and “hands”, requiring separate manual assistance for final diagnostic conclusion, which is not favorable for comprehensive Echo interpretation.
@@ 2.3 Towards domain-expertise agentic system
Recent days have witnessed growing interest in developing autonomous agentic systems, driven by the unparalleled tool-calling capability of MLLMs most recently [4, 34]. This promotes the development of agents to transcend the conversational nature of standard MLLMs by actively orchestrating external tools to acquire quantitative evidence and execute complex workflows [13, 32, 8, 11]. For instance,  [13] introduces a pipeline MedRAX, integrating specialized CXR tools with MLLMs to address complex chest X-ray analysis. In the more generalized CT and X‑ray domains, [32] enhances diagnostic performance by explicitly infusing MLLMs with knowledge from domain‑specialized vision models. Meanwhile, [8] presents a collaborative‑learning framework that adaptively incorporates pre-trained single-modal medical large models with small multimodal models to boost multimodal diagnosis. Despite these advances, we still lack a specialized agent for reliable Echo interpretation, since Echo is a unique imaging modality, characterized by its inherent multi‑view interdependency, complex anatomical geometry, and ambiguous structure boundaries [11]. These challenges necessitate a specialized synergy of Echo expertise with perceptual, operational, and reasoning capabilities, which most existing agents are not designed to fulfill.